OpenStudy (anonymous):

Eaton purchased a boat for $11,975. He made a down payment of $3,500. He applied for a five-year installment loan with an interest rate of 11.2%. What is the total cost of the boat after five years?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

ANYONE HELP?!?!

jimthompson5910 (jim_thompson5910):

completely stuck? or were you able to get somewhere?

OpenStudy (anonymous):

YAY JIM

jimthompson5910 (jim_thompson5910):

lol

OpenStudy (anonymous):

completely stuck, ive been at this question for like 20 mins, and im brain DEAD lol

jimthompson5910 (jim_thompson5910):

i see, ok well if the boat costs $11,975 (total cost + taxes etc etc) and he paid $3,500 as a down payment then how much is financed?

OpenStudy (anonymous):

i have like 9 questions left and im totally brain dead

OpenStudy (anonymous):

once again jim, im brain dead, might as well be a zombie

jimthompson5910 (jim_thompson5910):

you can tell me how much is financed at least

jimthompson5910 (jim_thompson5910):

how much does he need to borrow

OpenStudy (anonymous):

8475?

jimthompson5910 (jim_thompson5910):

so this is the principal of the loan

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

the term of the loan is 5 years

jimthompson5910 (jim_thompson5910):

interest rate (annual it looks like) is 11.25% and I'm going to assume interest is compounded monthly (that's usually how monthly payment deals are structured)

jimthompson5910 (jim_thompson5910):

11.2% I meant to say

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so to find the total cost of the loan, we need to find out his monthly payments first

OpenStudy (anonymous):

i got 14,508, but its not in options

jimthompson5910 (jim_thompson5910):

so we turn to this formula M = L[c(1+c)^n]/[(1+c)^n - 1] where M = monthly payment L = principal (amount loaned out) c = monthly interest rate n = number of terms (ie number of months)

OpenStudy (anonymous):

just tried it again and i got 14,798.46

OpenStudy (anonymous):

but thats wrong too

jimthompson5910 (jim_thompson5910):

in this case, M = unknown L = 8475 c = 0.112/12 = 0.0093333 n = 12*5 = 60 plug all this into the formula to get M = L[c(1+c)^n]/[(1+c)^n - 1] M = 8475*[0.0093333(1+0.0093333)^60]/[(1+0.0093333)^60 - 1] M = ??

OpenStudy (anonymous):

i got 182.26...

jimthompson5910 (jim_thompson5910):

hmm I'm getting M = 185.113303581369 ---> M = 185.11

jimthompson5910 (jim_thompson5910):

roundoff error most likely

OpenStudy (anonymous):

yeah, thats what im thinking

jimthompson5910 (jim_thompson5910):

anyways, that's how much he pays per month for 60 months

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

so in total, he will pay 60*185.11 = 11,106.60 dollars

OpenStudy (anonymous):

the choice that is closest is $11,106.81

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

I wonder if you're supposed to use tables somewhere

jimthompson5910 (jim_thompson5910):

instead of a calculator

OpenStudy (anonymous):

i have no clue, lol

OpenStudy (anonymous):

i dont think so

jimthompson5910 (jim_thompson5910):

maybe to calculate some piece (or all) of the monthly payment

OpenStudy (anonymous):

i dont even know how to do that then

jimthompson5910 (jim_thompson5910):

because usually, tables take complicated expressions and reduce them to simpler multipliers

jimthompson5910 (jim_thompson5910):

now that calculators have taken over, tables are a bit outdated, but some books may still use them

OpenStudy (anonymous):

yeah, i dont know haha

jimthompson5910 (jim_thompson5910):

hmm you'll have to ask your teacher about it

OpenStudy (anonymous):

so we cant do this one?

jimthompson5910 (jim_thompson5910):

might as well go with the closest thing then

OpenStudy (anonymous):

thats what im thinking

OpenStudy (anonymous):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it's not off by much tbh....but the fact that it's off is troubling

OpenStudy (anonymous):

i know right haha, my OCD is exploding

jimthompson5910 (jim_thompson5910):

21 cents is not a lot compared to the 11 thousand dollars, so it shouldn't matter lol

OpenStudy (anonymous):

true, thank bud

jimthompson5910 (jim_thompson5910):

np